OpenStudy (anonymous):

a car of mass 1200kg is brought to rest from a speed of 72km/h by a constant braking force of 3000 newtons. find how long the car takes to come to rest and the distance the car travels during this time?

OpenStudy (jamesj):

So first I'd convert the 72 km/h to SI units--meters/second. Now, having done that, what is the acceleration of a force of 3000 N? And hence how long does it take the car to stop?

OpenStudy (jamesj):

For the record, 72 km/hr = 20 m/s As F = ma, a force of -3000 N causes an deceleration on a mass of 1200 kg of a = -3000/1200 = -2.5 m/s^2 Hence at time t, the velocity of the car is v = 20 - 2.5t Now all you need to do to answer the first part is find the time t for which v = 0.